# Contextualized Messages Boost Graph Representations

### v2

@@ 3.1 Soft-isomorphic relational graph convolution network
For a graph representation learning problem, the relational message function $g$ may be modeled as a two-layered multi-layer perceptron (MLP), with an implied pseudometric, following the universal approximation theorem [16] to obtain the soft-isomorphic relational graph convolution network (SIR-GCN)
@@
$\mathcal{O}\left(\left|\mathcal{N}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{hidden}}+\left|\mathcal{N}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$ (11)
which is comparable to GAT. In practice, $\sigma$ may also be replaced with a deep MLP if modeling $g$ as a shallow two-layer MLP is infeasible. Moreover, since $\bm{h_{u}}$ is already encoded in the aggregation strategy, the combination strategy may simply be an activation function.
In essence, the proposed SIR-GCN is an instance of the MPNN framework where, unlike most MPNN instances in literature, the proposed model emphasizes the anisotropic and dynamic [7] transformation of the neighborhood features to obtain contextualized messages.
@@ 3.2 Soft-isomorphic graph readout function
Corollary 1 also shows that, for every graph $\mathcal{G}$, given a pseudometric distance $d_{\mathcal{G}}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{\mathcal{G}}$ on multisets of $\mathcal{H}$ defined in Eqn. 5, there exists a corresponding feature map $r_{\mathcal{G}}$ and soft-injective graph readout function $R_{\mathcal{G}}$ defined in Eqn. 7. While this result holds for every graph $\mathcal{G}$ independently, one may simply assume a single pseudometric $D$ with a corresponding soft-injective graph readout function $R$ for a set of graphs $\left\{\mathcal{G}_{d}\right\}_{d\in\mathcal{D}}$ from task $\mathcal{D}$. Nevertheless, one may opt to integrate additional information about the graph context and structure into the graph readout function $R$ to imitate having distinct $R_{\mathcal{G}}$ for every graph $\mathcal{G}$ and enhance its representational capability further. The virtual super node [14] may be used in this regard.
In practice, $R$ may also be modeled with an MLP, with an implied pseudometric, to obtain the soft-isomorphic graph readout function
$\bm{h_{\mathcal{G}}}=\sum_{v\in\mathcal{N}_{\mathcal{G}}}\textsc{MLP}_{R}\left(\bm{h_{v}}\right),$ (12)
where $\textsc{MLP}_{R}$ represents the corresponding feature map of $R$ and $\bm{h_{\mathcal{G}}}$ is the graph-level feature of graph $\mathcal{G}$.
@@ 4 Mathematical discussion
The mathematical relationship of SIR-GCN with GCN, GraphSAGE, GAT, GIN, and PNA are presented in this section to highlight the contribution. The relationship between SIR-GCN and the 1-WL test is also presented to contextualize the representational capability of the proposed model.
@@ 4.1 GCN and GraphSAGE
It may be shown that Corollary 1 holds up to a constant scale. Hence, the mean aggregation and symmetric mean aggregation (by extension) may be used in place of the sum aggregation. If one sets $\sigma$ as identity or $\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, $\bm{W_{R}}\bm{W_{K}}=\bm{W}$, and $\tilde{\mathcal{N}}(u)=\mathcal{N}(u)\cup\{u\}$, one obtains
@@
which recovers GCN and GraphSAGE with mean aggregation, respectively. Moreover, the sum aggregation may also be replaced with the max aggregation, albeit without theoretical justification, to recover GraphSAGE with max pooling. Thus, GCN and GraphSAGE may be viewed as instances of SIR-GCN. [^1] GraphSAGE with LSTM aggregation is not included in this discussion. The difference lies in the isotropic [12] nature of GCN and GraphSAGE and the use of non-linearities only in the combination strategy.
@@ 4.2 GAT
Moreover, in Brody et al. [7], the attention mechanism of GATv2 is modeled as an MLP given by
@@ 4.3 GIN
is equivalent to a GIN. Hence, SIR-GCN with residual connection encompasses GIN.
@@ 4.4 PNA
Furthermore, SIR-GCN and PNA approach the problem of uncountable node features $\mathcal{H}$ differently, with the former using only a single aggregator which holds theoretically for any number of neighbors. Nevertheless, both models highlight the significance of anisotropic message functions considering both the features of the query and key nodes. The key difference lies with PNA using a linear message function $m$ which is equivalent to a linear transformation of $\bm{h_{v}}$ with a different bias term for each $u$
@@
When using mean, max, or min aggregators, the query node $u$ only contributes to the aggregated neighborhood features through the bias term $\bm{b_{u}}$. Meanwhile, when using normalized moment aggregators, the query node $u$ no longer contributes to the aggregated neighborhood features. This potentially limits the expressivity of PNA since, in contrast to SIR-GCN and as suggested by Brody et al. [7], messages are only linearly and statically [7] transformed by the features of the query node.
@@
Overall, SIR-GCN is demonstrated to generalize four prominent GNNs in literature – GCN, GraphSAGE, GAT, and GIN – and is thus at least as expressive as they are. Notably, SIR-GCN offers flexibility in two dimensions of GNNs: aggregation strategy and message transformation. It emphasizes summed aggregation with anisotropic and dynamic (i.e. contextualized) message transformation, making it well-suited for heterophilous tasks [8], but remains adaptable to alternative configurations.
In addition, SIR-GCN distinguishes itself from PNA by employing only a single aggregator, which theoretically holds for graphs of arbitrary sizes, thus reducing computational complexity. Nevertheless, its expressivity is maintained through contextualized messages, allowing it to inherit the representational capability of the 1-WL test.
@@
Figure 3: SIR-GCN expressivity.
@@ 5 Experiments
Experiments on synthetic and benchmark datasets in node and graph property prediction tasks are conducted to highlight the expressivity of SIR-GCN. To ensure fairness, models not employing advanced architectural design or manually crafted features using domain knowledge are used as primary comparisons.
@@ 5.1 Synthetic datasets
DictionaryLookup [7] consists of bipartite graphs with $2n$ nodes – $n$ key nodes each with an attribute and value and $n$ query nodes each with an attribute. The task is to predict the value of query nodes by matching their attribute with the key nodes. A sample instance is provided in Fig. 4.
@@
$A,1$$B,2$$C,3$$D,4$$A,*$$B,*$$C,*$$D,*$
Figure 4: DictionaryLookup.
$A$$B$$A$$B$
Figure 5: GraphHeterophily.
Table 1 presents the mean and standard deviation of the test accuracy for SIR-GCN, GCN, GraphSAGE, GATv2, and GIN across different values of $n$. SIR-GCN and GATv2 achieve perfect accuracy attributed to their anisotropic nature. However, it is observed that GATv2 suffers from performance degradation in some trials. Meanwhile, the other models fail to predict the value of query nodes even for the training graphs due to their isotropic nature. The results underscore the utility of an attentional/relational mechanism in capturing the relationship between the query and key nodes.
@@
GraphHeterophily is an original synthetic dataset. It consists of random directed graphs with each node assigned one of $c$ classes. The task is to count the number of directed edges connecting two nodes with distinct class labels within each graph. A sample instance is provided in Fig. 5.
@@
Table 2 presents the mean and standard deviation of the test mean squared error (MSE) for SIR-GCN, GCN, GraphSAGE, GATv2, and GIN across different values of $c$. SIR-GCN achieves an MSE loss of nearly 0 attributed to its anisotropic nature and sum aggregation. In contrast, GCN, GraphSAGE, and GATv2 obtained large MSE losses due to their mean or max aggregation which fails to preserve the graph structure as noted by Xu et al. [47]. Meanwhile, GIN successfully retains the graph structure but fails to learn the relationship between the labels of the query (center) node and key (neighboring) nodes. The results illustrate the utility of anisotropic models even in graph property prediction tasks with countable node features.
@@ 5.2 Benchmark datasets
Benchmarking GNNs [12] is a collection of benchmark datasets consisting of diverse mathematical and real-world graphs across various GNN tasks. In particular, the WikiCS, PATTERN, and CLUSTER datasets fall under node property prediction tasks while the MNIST, CIFAR10, and ZINC datasets fall under graph property prediction tasks. Furthermore, the WikiCS, MNIST, and CIFAR10 datasets have uncountable node features while the remaining datasets have countable node features. The performance metric for ZINC is the mean absolute error (MAE) while the performance metric of the remaining datasets is accuracy. Dwivedi et al. [12] provides more information regarding the individual datasets.
@@
Table 3 presents the mean and standard deviation of the test performance for SIR-GCN and comparable GNN models across the six benchmarks where the experimental set-up follows that of Dwivedi et al. [12] to ensure fairness. The results show that SIR-GCN consistently outperforms popular GNNs in literature. Notably, SIR-GCN also outperforms both PNA [11] and efficient graph convolution (EGC-M) [39] which use multiple aggregators. This highlights the significance of non-linear and contextualized messages in enhancing the expressivity of GNNs, complementing the discussion in the previous section.
@@
ogbn-arxiv [19] is a benchmark dataset representing the citation network between all Computer Science (CS) arXiv papers indexed by Microsoft academic graph [42]. Each node represents an arXiv paper and a directed edge represents a citation. The task is to classify each paper, based on its title and abstract, into the 40 subject areas of arXiv CS papers.
@@
Model GIANT-XRT [10] BoT [45] C&S [20] Others Accuracy Parameters
@@
Table 4 presents the mean and standard deviation of the test accuracy for SIR-GCN and other models in literature. The tricks used and the number of parameters are also presented for completeness. The results show that SIR-GCN, utilizing only a single GNN layer, outperforms comparable models in predicting the subject area of the papers. Unsurprisingly, however, SIR-GCN performs poorly when compared against complex frameworks utilizing more tricks such as the dynamic evolving initial residual GAT (DRGAT) [1], reversible GAT (RevGAT) [26], and adaptive graph diffusion network (AGDN) [38], albeit with only a small performance difference.
@@
ogbg-molhiv [19] is another benchmark dataset where each graph represents a molecule with nodes representing atoms and edges representing chemical bonds. Node features contain information regarding the atom while edge features contain information regarding the chemical bond. The task is to predict whether or not the molecules inhibit HIV replication.
@@
Model GraphNorm [9] VirtualNode [14] Others ROC-AUC Parameters
@@
GSN [6] 0.7799 $\pm$ 0.0100 3,338,701
GSAT [31] 0.8067 $\pm$ 0.0950 249,602
CIN [4] 0.8094 $\pm$ 0.0057 239,745
Table 5 presents the mean and standard deviation of the test ROC-AUC for SIR-GCN and other models in literature. The tricks used and the number of parameters are also presented for completeness. The results show that with only a single GNN layer, SIR-GCN outperforms established models in predicting molecules inhibiting HIV replication, highlighting its expressivity. Nevertheless, SIR-GCN fails to compete with more complex models such as the graph stochastic attention (GSAT) [31] and models incorporating domain knowledge such as the cell isomorphism network (CIN) [4].
@@ 6 Conclusion
Overall, the paper provides a novel perspective for creating a powerful GNN when the space of node features is uncountable. The central idea is to use implicit pseudometric distances to create soft-injective functions such that distinct inputs may produce similar outputs only if the distance between the inputs is sufficiently small on some representation. This concept is demonstrated at all levels from node features to graph-level features. From the results, a novel SIR-GCN is proposed and shown to generalize classical GNN methodologies. The expressivity of SIR-GCN is then empirically demonstrated with synthetic and benchmark datasets, highlighting its relative superiority in outperforming comparable models. Future studies may explore incorporating SIR-GCN into existing frameworks such as using multiple aggregators [11], reversible GNNs [26], and adaptive graph diffusion networks [38] to achieve state-of-the-art performance.
Acknowledgments and Disclosure of Funding

### v3

@@ 3.1 Soft-isomorphic relational graph convolution network
For a graph representation learning problem, one may directly model the anisotropic and dynamic relational message function $g$ as a two-layer multi-layer perceptron (MLP), with implicitly learned pseudometrics, following the universal approximation theorem (Hornik et al., 1989) to obtain the soft-isomorphic relational graph convolution network (SIR-GCN)
@@
$\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{hidden}}+\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$ (11)
with computational efficiency achieved by the application of only an activation function along edges, making it comparable to GAT. Nevertheless, $\sigma$ may also be replaced with a deep MLP in practice if modeling $g$ as a shallow two-layer MLP becomes infeasible.
In essence, the proposed SIR-GCN is a simple, interpretable, and computationally efficient instance of the MPNN framework. Moreover, in contrast to other MPNN instances in literature, the proposed model emphasizes the anisotropic and dynamic transformation of neighborhood features to obtain contextualized messages.
@@ 3.2 Soft-isomorphic graph readout function
Corollary 1 also shows that, for each graph $\mathcal{G}$, given a pseudometric distance $d_{\mathcal{G}}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{\mathcal{G}}$ on multisets of $\mathcal{H}$ defined in Eqn. 5, there exists a corresponding feature map $r_{\mathcal{G}}$ and soft-injective graph readout function $R_{\mathcal{G}}$ defined in Eqn. 7. While this result holds for each graph $\mathcal{G}$ independently, one may simply consider a single $D$, $d$, $r$, and $R$ for every graph $\left\{\mathcal{G}_{d}\right\}_{d\in\mathcal{D}}$ under task $\mathcal{D}$. Nevertheless, the graph context and structure may also be integrated into $D$, $d$, $r$, and $R$, through a virtual super node (Gilmer et al., 2017) for instance, to imitate having distinct $D_{\mathcal{G}}$, $d_{\mathcal{G}}$, $r_{\mathcal{G}}$, and $R_{\mathcal{G}}$ for each graph $\mathcal{G}$ and to further enhance its representational capability.
Similarly, for a graph representation learning problem, $r$ may also be directly modeled as an MLP, with implicitly learned pseudometrics, to obtain the soft-isomorphic graph readout function
$\bm{h_{\mathcal{G}}}=\sum_{v\in\mathcal{V}_{\mathcal{G}}}\textsc{MLP}_{R}\left(\bm{h_{v}}\right),$ (12)
where $\textsc{MLP}_{R}$ corresponds to $r$ and $\bm{h_{\mathcal{G}}}$ is the graph-level feature of graph $\mathcal{G}$.
@@ 4 Mathematical discussion
The mathematical relationship of SIR-GCN with GCN, GraphSAGE, GAT, GIN, and PNA are presented in this section to highlight the novelty and contribution. While activation functions and MLPs applied after each GNN layer play a significant role in the overall performance, the discussions only focus on the message-passing operation that defines GNNs. The relationship between SIR-GCN and the 1-WL test is also presented to contextualize the representational capability of the former.
@@ 4.1 GCN and GraphSAGE
It may be shown that Corollary 1 holds up to a constant scale. Hence, the mean aggregation and symmetric mean aggregation, by extension, may be used in place of the sum aggregation. If one sets $\sigma$ as identity or $\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, $\bm{W_{R}}\bm{W_{K}}=\bm{W}$, and $\tilde{\mathcal{N}}(u)=\mathcal{N}(u)\cup\{u\}$, one obtains
@@
which recovers GCN and GraphSAGE with mean aggregation, respectively. Moreover, the sum aggregation may also be replaced with the max aggregation, albeit without theoretical justification, to recover GraphSAGE with max pooling. Thus, GCN and GraphSAGE may be viewed as instances of SIR-GCN. [^2] GraphSAGE with LSTM aggregation is not included in this discussion. The difference lies in the isotropic (Dwivedi et al., 2023) nature (i.e., a function of only the features of the key nodes) of GCN and GraphSAGE and the use of non-linearities only in the combination strategy.
@@ 4.2 GAT
Moreover, in Brody et al. (2021), the attention mechanism of GATv2 is modeled as an MLP given by
@@ 4.3 GIN
is equivalent to GIN. Hence, SIR-GCN with residual connection encompasses GIN.
@@ 4.4 PNA
Furthermore, while SIR-GCN and PNA approach the problem of uncountable node features differently, both models highlight the significance of anisotropic message functions considering both the features of the query and key nodes. The key difference lies with PNA using a static (Brody et al., 2021) or linear message function $m$ which translates to
@@
As a result, the influence of the query node on the aggregated neighborhood features is limited. For instance, when using mean, max, or min aggregators, the influence of the query node $u$ is restricted to the bias term $\bm{b_{u}}$. Moreover, with normalized moment aggregators, the bias term is effectively canceled out during the normalization process, further reducing the influence of the query node. Hence, PNA does not fully leverage its anisotropic nature, limiting its expressivity. In contrast, the dynamic nature of SIR-GCN allows for the non-linear embedding of the features of the query node $\bm{h_{u}}$ within the aggregated neighborhood features, thereby fully leveraging its anisotropic nature.
@@
4.6 SIR-GCN
Overall, SIR-GCN offers flexibility in two key dimensions of GNNs: aggregation strategy and message transformation. Consequently, it generalizes four prominent GNNs in literature—GCN, GraphSAGE, GAT, and GIN—ensuring that it is at least as expressive as these models. Notably, SIR-GCN sets itself apart from other GNNs as the first MPNN instance to incorporate both anisotropic and dynamic (i.e., contextualized) messages within the MPNN framework, making it well-suited for heterophilous tasks (Bronstein et al., 2021) while remaining adaptable to homophilous tasks.
@@ 4.6 SIR-GCN
Figure 3: SIR-GCN encompasses GCN, GraphSAGE, GAT, and GIN.
In addition, SIR-GCN distinguishes itself from PNA by employing only a single aggregator that theoretically holds for graphs of arbitrary sizes, thus reducing computational complexity. Nevertheless, its expressivity is maintained through contextualized messages, allowing it to inherit the representational capability of the 1-WL test.
@@ 5 Experiments
Experiments on synthetic and benchmark datasets in node and graph property prediction tasks are conducted to highlight the expressivity of SIR-GCN. To ensure fair evaluation, models not employing complex architectural design or manually crafted features using domain knowledge are used as primary comparisons.
@@ 5.1 Synthetic datasets
DictionaryLookup (Brody et al., 2021) consists of bipartite graphs with $2n$ $n$ key nodes each with an attribute and value and $n$ query nodes each with an attribute. The task is to predict the value of query nodes by matching their attribute with the key nodes.
$A,1$$B,2$$C,3$$D,4$$A,*$$B,*$$C,*$$D,*$
Figure 4: DictionaryLookup.
$A$$B$$A$$B$
Figure 5: GraphHeterophily.
@@
Table 1 presents the mean and standard deviation of the test accuracy for SIR-GCN, GCN, GraphSAGE, GATv2, GIN, and PNA across different values of $n$. SIR-GCN and GATv2 achieve perfect accuracy attributed to their anisotropic and dynamic nature. However, it is observed that GATv2 suffers from performance degradation in some trials. Meanwhile, the other models fail to predict the value of query nodes even for the training graphs due to their isotropic and/or static nature. The results underscore the utility of a dynamic attentional or relational mechanism in capturing the relationship between the query and key nodes.
@@
GraphHeterophily is an original synthetic dataset. It consists of random directed graphs with each node assigned one of $c$ classes. The task is to count the number of directed edges connecting two nodes with distinct class labels within each graph.
@@
Table 2 presents the mean and standard deviation of the test mean squared error (MSE) for SIR-GCN, GCN, GraphSAGE, GATv2, GIN, and PNA across different values of $c$. SIR-GCN achieves an MSE loss of nearly zero attributed to its anisotropic and dynamic nature and sum aggregation. In contrast, GCN, GraphSAGE, and GATv2 obtained large MSE losses due to their mean or max aggregation which fails to preserve the graph structure as noted by Xu et al. (2018a). Meanwhile, GIN and PNA successfully retain the graph structure but fail to learn the relationship between the labels of the query node and key nodes due to their static nature. The results illustrate the utility of anisotropic and dynamic models using sum aggregation even with countable node features.
@@ 5.2 Benchmark datasets
Benchmarking GNNs (Dwivedi et al., 2023) is a collection of benchmark datasets consisting of diverse mathematical and real-world graphs across various GNN tasks. In particular, the WikiCS, PATTERN, and CLUSTER datasets fall under node property prediction tasks while the MNIST, CIFAR10, and ZINC datasets fall under graph property prediction tasks. Furthermore, the WikiCS, MNIST, and CIFAR10 datasets have uncountable node features while the remaining datasets have countable node features. The performance metric for ZINC is the mean absolute error (MAE) while the performance metric of the remaining datasets is accuracy. Dwivedi et al. (2023) provides more information regarding the individual datasets.
@@
Table 3 presents the mean and standard deviation of the test performance for SIR-GCN and comparable GNN models across the six benchmarks where the experimental set-up follows that of Dwivedi et al. (2023) to ensure fair evaluation. The results show that SIR-GCN consistently outperforms popular GNNs in literature. Notably, SIR-GCN also outperforms both PNA (Corso et al., 2020) and efficient graph convolution (EGC-M) (Tailor et al., 2021) which use multiple aggregators. This highlights the significance of contextualized messages in enhancing the expressivity of GNNs, complementing the discussion in the previous section.
@@
ogbn-arxiv (Hu et al., 2020) is a benchmark dataset representing the citation network between all Computer Science (CS) arXiv papers indexed by Microsoft academic graph (Wang et al., 2020). Each node represents an arXiv paper and a directed edge represents a citation. The task is to classify each paper, based on its title and abstract, into the 40 subject areas of arXiv CS papers.
@@
Model GIANT-XRT (Chien et al., 2021) BoT (Wang et al., 2021) C&S (Huang et al., 2020) Others Accuracy Parameters
@@
Table 4 presents the mean and standard deviation of the test accuracy for SIR-GCN and other models in literature. The tricks used and the number of parameters are also presented for completeness. The results show that SIR-GCN, utilizing only a single GNN layer, outperforms comparable models in predicting the subject area of the papers. As expected, however, SIR-GCN fails to compete with complex frameworks utilizing more tricks such as the reversible GAT (RevGAT) (Li et al., 2021) and the adaptive graph diffusion network (AGDN) (Sun et al., 2020), both of which build upon GAT by employing grouped reversible residual connections and adaptive graph diffusion, respectively. Nevertheless, SIR-GCN achieves performance close to that of the complex GNN frameworks mentioned, showcasing an effective balance between complexity and expressivity.
@@
ogbg-molhiv (Hu et al., 2020) is another benchmark dataset where each graph represents a molecule with nodes representing atoms and edges representing chemical bonds. Node features contain information regarding the atom while edge features contain information regarding the chemical bond. The task is to predict whether or not the molecules inhibit HIV replication.
@@
Model GraphNorm (Cai et al., 2021) VirtualNode (Gilmer et al., 2017) Others ROC-AUC Parameters
@@
GSN 0.7799 $\pm$ 0.0100 3,338,701
GSAT 0.8067 $\pm$ 0.0950 249,602
CIN 0.8094 $\pm$ 0.0057 239,745
Table 5 presents the mean and standard deviation of the test area under the receiver operating characteristic curve (ROC-AUC) for SIR-GCN and other models in literature. The tricks used and the number of parameters are also presented for completeness. The results show that with only a single GNN layer, SIR-GCN outperforms comparable models in predicting molecules inhibiting HIV replication, highlighting its expressivity. Given the simplicity of SIR-GCN, it is expected to exhibit lower performance compared to complex models such as the graph stochastic attention (GSAT) (Miao et al., 2022), which builds upon PNA by leveraging the information bottleneck principle, and the cell isomorphism network (CIN) (Bodnar et al., 2021), which is a hierarchical message-passing framework utilizing the topological features of graphs. Despite its simple design, the expressivity of SIR-GCN is evident in its close performance to that of complex GNN frameworks.
@@ 6 Conclusion
In summary, the paper provides a novel perspective for creating a powerful GNN across all levels when the space of node features is uncountable. The central idea is to use pseudometric distances to create soft-injective functions such that distinct inputs may produce similar outputs if and only if the distance between inputs is sufficiently small on some representation. From the results, the SIR-GCN is proposed as the first MPNN instance to emphasize contextualized message transformation, setting it apart from other GNNs. Furthermore, the model is shown to generalize classical GNN methodologies. Despite its simple design, empirical results underscore SIR-GCN as the best-performing MPNN instance that effectively balances model complexity and expressivity. The paper thus contributes to GNN literature by theoretically and empirically demonstrating the necessity of both anisotropic and dynamic messages to enhance GNN expressivity. Future works may consider incorporating SIR-GCN into complex frameworks, such as grouped reversible residual connections (Li et al., 2021), adaptive graph diffusion (Sun et al., 2020), graph stochastic attention (Miao et al., 2022), and hierarchical message-passing (Bodnar et al., 2021), to address the limitations inherent in the MPNN framework and to develop more expressive GNN architectures.
Acknowledgments